### v2

# xGen-MM (BLIP-3): A Family of Open Large Multimodal Models

Abstract
This report introduces xGen-MM (also known as BLIP-3), a framework for developing Large Multimodal Models (LMMs). The framework comprises meticulously curated datasets, a training recipe, model architectures, and a resulting suite of LMMs. xGen-MM, short for xGen-MultiModal, expands the Salesforce xGen initiative on foundation AI models. Our models undergo rigorous evaluation across a range of tasks, including both single and multi-image benchmarks. Our pre-trained base model exhibits strong in-context learning capabilities and the instruction-tuned model demonstrates competitive performance among open-source LMMs with similar model sizes. In addition, we introduce a safety-tuned model with DPO, aiming to mitigate harmful behaviors such as hallucinations and improve safety. We open-source our models, curated large-scale datasets, and our fine-tuning codebase to facilitate further advancements in LMM research. Associated resources will be available on our project page above. Refer to caption Refer to caption (a) BLIP-2 framework (b) xGen-MM (BLIP-3) framework Figure 1: We introduce xGen-MM (BLIP-3), a framework (b) for developing Large Multimodal Models (LMMs). Our framework improves upon BLIP-2 (a) [1] by (1) increasing the richness, scale, and diversity of training data, (2) replacing the Q-Former layers with a more scalable vision token sampler, and (3) simplifying the training process via the unification of the training objectives to a single loss at every training stage. The resulting suite of LMMs can perform various visual language tasks and achieve competitive performance across benchmarks.
1 Introduction
Large Multimodal Models (LMMs) have attracted significant attention with their potential applications and emergent capabilities. Recent advancements in both proprietary models [2, 3, 4, 5] and open-source LMMs [6, 1, 7, 8, 9, 10, 11] highlight the rapid progress and growing interest in this field. However, despite these advancements, there is still a gap between open-source models and proprietary ones in terms of access to open weights, training recipes, and curated datasets. Such limitations hinder the open-source communities from replicating, understanding, and improving LMMs.
Recent works have demonstrated that large-scale and high-quality data are essential for training robust LMMs [8, 9, 10, 11, 12]. BLIP-2 [1] was one of the pioneering efforts in exploring LMMs, which leveraged synthetic data to achieve impressive results at the time (Figure 1 (a)). However, the data used in BLIP-2 lacks the scale, quality, and diversity required to reach competitive performance compared to more modern LMMs nowadays. In addition, BLIP-2 employs an intricate Q-Former [1] architecture to bridge the vision and language modalities, coupled with a suite of complex training objectives (ITM, ITC, and ITG losses), both of which pose obstacles for larger-scale training. Moreover, BLIP-2 supports only single-image input, whereas interleaved multimodal data formats are the most natural form of multimodal data [13].
In response to these challenges, we introduce xGen-MM (BLIP-3) (Figure 1 (b)), a new framework designed to scale up LMM training by utilizing an ensemble of multimodal interleaved datasets, curated caption datasets, and other publicly available datasets [14, 15, 16, 17]. xGen-MM, short for xGen-MultiModal, further expands our previous generative AI initiatives and corresponding foundation models for text xGen [18], code generation codeGen [19, 20], function calling APIGen [21], among others. In xGen-MM (BLIP-3), as illustrated in Figure 2, we streamline the model architecture by replacing the Q-Former [1] with a more scalable vision token sampler (specifically, a perceiver resampler [22]) and simplifying the training objectives to focus solely on the auto-regressive loss of text tokens in a multimodal context. Our primary focus is on dataset curation and scaling up the training data. Recently, our BLIP-3 team introduced two large-scale, high-quality datasets: MINT-1T [12], a trillion-token scale interleaved dataset; and BLIP3-KALE, a knowledge-augmented high-quality dense captions dataset. In this technical report, we introduce two additional specialized datasets: BLIP3-OCR-200M, a large-scale dataset with dense OCR annotations; and BLIP3-GROUNDING-50M, a large-scale visual grounding dataset.
In addition to these datasets, we are committed to open-sourcing the series of models developed in this work, including the base, instruction-tuned, and DPO models. Along with the model release, we also provide code for easy fine-tuning of our base model on custom datasets. By making these resources publicly available, we aim to make LMM research and development more accessible to the community, and we encourage researchers and practitioners to use our models and datasets to understand and further explore the potential and emergent abilities of LMMs.
Figure 2: Overview of the xGen-MM (BLIP-3) framework. Free-form interleaved images and texts from the ensembled interleaved and caption datasets are input into the framework, with each modality undergoing a separate tokenization process to be fed into the pre-trained LLM in natural order. A standard auto-regressive loss is then applied to the text tokens. The Vision Transformer is kept frozen during training, while all other parameters, including the token sampler and the pre-trained LLM, are trained.
2 Related Work
Recent advancements in Large Multimodal Models (LMMs) have explored two main architectural approaches: the cross-attention style [22, 23] and the decoder-only style. The cross-attention approach, exemplified by models like Flamingo [22, 23] and Llama 3.1 [5], integrates vision and language modalities through a complex attention mechanism to enable deep multimodal understanding. Another approach is the decoder-only architecture [1, 7, 8, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], which we adopt in xGen-MM (BLIP-3), offers a more streamlined solution. This approach connects pre-trained language models to visual inputs using lightweight connectors, simplifying the integration process while maintaining robust multimodal capabilities. The effectiveness of this architecture is evident in models such as MM1 [9], VILA [10], LLaVA [8], phi3-vision [37], and Otter [38].
Training methodologies for LMMs typically follow one of the two strategies. The first one uses a light pre-training procedure and heavily relies on visual instruction tuning, as seen in the LLaVA series [8, 29]. Extensive research has been conducted on creating effective instruction-tuning data for a variety of tasks [32, 39, 40, 41, 42, 43]. The second strategy involves extensive pre-training on large-scale, diverse datasets, followed by visual instruction fine-tuning. This approach, employed by models like MM1 and Idefics2 [11], infuses broad knowledge into the model, which is then fine-tuned to align with human-like interactions and safety standards. While MM1 [9] provides extensive ablations and studies on the recipes aimed at improving LMMs, it releases limited resources for practitioners to adopt the model (e.g., MM1 models and datasets are close-sourced). Idefics2 [11] is a more recent open-source work that open-sources a suite of models along with detailed training strategies and data recipes, but Idefics2 mostly resues existing datasets in their experiments without contributing new ones.
In this work, we present xGen-MM (BLIP-3). Unlike previous works, xGen-MM (BLIP-3) is an open-source family of a series of models, data recipes, fine-tuning code, and two large-scale foundational multimodal datasets, which we hope will enable and advance future research in this area.
3 Model Architecture
Architecture Overview.
As illustrated in Figure 2, the xGen-MM (BLIP-3) framework adopts an architecture consisting of a ViT [44, 45], a vision token sampler (perceiver resampler [22]) to downsample the image embeddings, and a pre-trained Large Language Model (phi3-mini [37]). The input to the model can be free-form multimodal interleaved texts and vision tokens from the diverse multimodal data sources we ensemble.
Any-Resolution Vision Token Sampling.
As proved effective in recent LMMs [46, 47, 48], we adopt a dynamic high-resolution (i.e., "any-resolution") image encoding strategy at the fine-tuning and post-training stages. We enable higher-resolution image understanding with image patch-wise encoding. The patch-wise encoding preserves the resolution of the original image as much as possible by splitting a single image into multiple patches and encoding them separately. Following the previous convention, we concatenate the encoded image patches with a downsized original image that provides global information.
In the VL connector, we use a perceiver resampler to downsample the vision tokens. With any-resolution image encoding, we perform the downsampling for each image patch (including the downsized original image) independently. The downsampler vision tokens are then concatenated together and sent to the LLM. With the downsampling in our VL connector, we can reduce the sequence length of vision tokens by a factor of five or more depending on the number of query tokens in the perceiver resampler. We provide ablation studies on different token sampling strategies in Section 7.2.
4 Training
Pre-training.
The pre-training objective is to predict the next text token across the dataset mixture we pre-train on. Overall, the resulting base model xGen-MM-Phi3-mini-base-r is pre-trained for about 100 billion multimodal tokens from the ensembled dataset, and our pre-training resolution is 384x384 pixels, which aligns with SigLIP [45].
Supervised Fine-tuning (SFT).
We further fine-tune our pre-trained models on instruction-following examples to make them better understand and follow user queries. At the fine-tuning stage, we use a collection of publically available instruction-following datasets [11, 29, 49]. We adopt the any-resolution vision token sampling strategy to allow a better understanding of images of higher resolutions such as text-rich document data. We introduce the technical details for the fine-tuning stage in the following sections.
Interleaved Multi-Image Supervised Fine-tuning.
We conduct a second-stage fine-tuning on the instruction fine-tuned model on a mixture of multi-image and single-image instructions-following samples. The goal for this second-stage fine-tuning is to enhance the model’s ability to comprehend interleaved image-text input, which is helpful for multimodal in-context learning, multi-image question answering, and many more practical use cases. For the multi-image fine-tuning, we also adopt the any-resolution vision token sampling strategy same as in the previous SFT stage.
Post-training.
Finally, we perform two stages of post-training to improve the model’s helpfulness while mitigating harmful qualities such as hallucination and toxicity. We first perform direct preference optimization (DPO [50]) to improve the model’s helpfulness and visual faithfulness. We then perform safety fine-tuning to improve the model’s harmlessness. We quantitatively demonstrate Pareto gains in model harmlessness and helpfulness after our post-training.
5 Data
5.1 Pre-training Data Recipe
As indicated in Figure 3, in xGen-MM (BLIP-3), we pre-train on an ensemble of diverse multimodal datasets with the indicated sampling ratios.
Figure 3: Overview of xGen-MM (BLIP-3) Pre-training Datasets.
Interleaved Dataset Mixture.
We combine MINT-1T (including its HTML, PDF, and ArXiv subsets) with OBELICS (HTML only) to create a more diverse and comprehensive dataset mixture that covers a broader range of domains.
  1. 1.
    MINT-1T [12] is a trillion token scale multimodal interleaved dataset, containing data sources from HTML, PDF, and ArXiv. As evidenced by MM1 [9] and Idefics2 [11], such multimodal interleaved datasets are essential for scaling up large multimodal model training and enabling fundamental capabilities like multimodal in-context learning. Notably, different from the OBELICS [11], MINT-1T has three subsets from different sources: the HTML subset, the PDF subset, and the ArXiv subset. In xGen-MM (BLIP-3), these three subsets are mixed in a 7:5:1 ratio.
  2. 2.
    OBELICS [11] is another large-scale multimodal interleaved dataset constructed from HTML documents solely. It differs slightly in domain coverage from MINT-1T due to the specific preprocessing steps adopted.
Caption Dataset Mixture.
We integrate a diverse range of caption datasets, with the specifics outlined in the following details.
  1. 1.
    BLIP3-KALE is a large-scale curated high-quality caption dataset. Details will be discussed in another paper, and the dataset will be made public very soon.
  2. 2.
    BLIP3-OCR-200M is a curated large-scale OCR dataset to address the limitations of current large multimodal models in handling text-rich images like documents and charts, as traditional image-text datasets often lack adequate OCR annotations. To enhance text comprehension abilities, we use the OCR engine PaddleOCR [51] to annotate images with OCR-specific annotations. Overall, we curate a dataset of 200 million high-resolution images from Datacomp-1B[17]. For each image, we create captions with OCR data by identifying and extracting textual elements using the off-the-shelf OCR engine [51]. Text segments in a caption like "… text …" are modified to include OCR information as "… text ( ocr_info ) …", where ocr_info contains bounding box coordinates for the extracted text, specifying its exact position within the image in the format "<bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>". We have multiple granularities of OCR information, including with and without bounding box data. In our work, we only utilize textual information without bounding box data, which has proven to be effective. Note that, in xGen-MM (BLIP-3), we preprocess the captions to remove filler texts such as “the text”, which we find improves OCR-related benchmark performance. We hypothesize that this is because such filler text is relatively easy to predict, potentially diluting the loss associated with OCR-relevant tokens. Nonetheless, incorporating bounding box information could further enhance performance, and we encourage researchers in the community to explore this potential.
  3. 3.
    BLIP3-GROUNDING-50M is a curated large-scale grounding dataset to enhance the ability to ground semantic concepts in visual features, which is crucial for tasks like object detection, semantic segmentation, and understanding referring expressions [52] (e.g., "the object to the left of the dog"). We curate a dataset of 50 million images from Datacomp-1B [17]. For each image, we identify objects and their location information using one of the state-of-the-art open-world image tagging [53] and object detection models [54]. Objects mentioned in a caption like "… object …" are modified to include grounding information as " … object ( grounding_info ) …", where grounding_info contains bounding box information in one of three formats, each capturing a different granularity of localization: (1) <bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>, (2) "starts at $(x_{1},y_{1})$ and extends up to $(x_{2},y_{2})$", or (3) "top-left corner of the image".
  4. 4.
    Other Public Datasets Mixture: We also include other publicly available datasets such as uncurated Datacomp-1B [17] image-text pairs, CC12M [14], CC3M [14], VG [15], and SBU [16].
Figure 4: Samples from BLIP3-OCR-200M. Six levels of OCR information granularity are extracted, with and without bounding box data. Note that in xGen-MM (BLIP-3), OCR-related captions are preprocessed to remove filler phrases like ’the text,’ resulting in improved OCR benchmark performance.
Figure 5: Samples from BLIP3-GROUNDING-50M. We introduce a large-scale dataset of images and corresponding captions containing localization information about objects. Furthermore, we release the associated object bounding box data to facilitate the creation of captions with custom templates.
5.2 Supervised Fine-tuning Data Recipe
The datasets used in the fine-tuning stage are from public datasets of different domains [29, 11, 49]. We sample data with various domain focuses including multi-modal conversation [29], image captioning [55, 56], visual question answering [57, 58, 59, 60], chart/document understanding [61, 62, 63, 64], science and math [65, 66]. In addition to the multi-modal image-text data, we also mix in pure text instruction following data  [67, 68] during visual instruction tuning. Ultimately, we collect a mixture of 1 million publically available instruction-tuning samples, on which we fine-tune our model for one epoch.
The multi-image instruction tuning stage starts with a model fine-tuned on single-image samples. We use a mixture of public multi-image/interleaved image-text instruction data [69, 70]. To prevent the model from deteriorating on single-image capabilities, we reuse a subset of single-image datasets used in the previous fine-tuning stage and mix them into the multi-image training data.
5.3 Post-training Data Recipe
Improving Truthfulness by Direct Preference Optimization.
We employ VLFeedback [71], a synthetically annotated multimodal preference dataset that uses off-the-shelf VLMs to generate responses to a diverse mix of multimodal instructions that are then scored by GPT4-V [2] along three axes – helpfulness, visual faithfulness, and ethics. The dataset contains 80k such instructions from which we construct preference data by marking as preferred (and dispreferred) the response with the highest (and lowest) average score across models and filtering out examples with low-scoring preferred responses. We thus generate 62.6k preference examples.
We perform 1 epoch of DPO on the combined preference dataset while updating a subset (2.5%) of LLM backbone weights using low-rank adaptation (LoRA [72]). Also, following recent work [50], we generate an additional set of responses that capture the model’s intrinsic hallucinations, by performing a second step of DPO per-iteration against the models’ output to a noised version of the input image and original query, which we treat as an additional dispreferred response.
Improving Harmlessness by Safety Fine-tuning.
Next, we perform 3 epochs of safety fine-tuning on the train split of the VLGuard [73] dataset, which contains 2k examples of unsafe images and instructions. VLGuard comprises two types of unsafe examples: (1) objectionable images paired with safe instructions and a desirable abstention response, and (2) safe images paired with two types of instruction-response pairs, one safe and another unsafe. The dataset consists of unsafe examples belonging to various subcategories including privacy-violating, risky/sensitive topics (such as politics, sex, and violence), deception, and discrimination. Following the original work, we randomly sample 5k additional examples from the instruction fine-tuning dataset to retain the model’s helpfulness without exaggerating its safety behavior. As before, we update a subset (2.5%) of LLM backbone weights using low-rank adaptation.
6 Experiments
6.1 Pre-training
Few-shot Evaluation.
After the pre-training stage, we evaluate our pre-trained model on classic captioning and VQA tasks, in comparison with previous models that support few-shot learning multi-modal evaluation. We present zero-shot and few-shot (4- and 8-shots) results, as shown in Table 1. Overall, our model achieves competitive multimodal in-context learning performance with comparable-sized LMMs  [^1] For few-shot evaluation, the zero-shot results are used mainly as a reference for their corresponding few-shot numbers, and they can be sensitive to prompts. As also mentioned in [9], they are mostly indicative of how well the pre-training distribution matches the associated evaluation task format. In pre-training evaluation, we mainly care for few-shot performance, which is robust to prompt templates.. For the OCR tasks (TextCaps and TextVQA) and VQA-v2, it significantly outperforms MM1-3B and even larger models such as Idefics-9B [13] and MM1-7B [9]. On all benchmarks, increasing the number of shots can improve the performance, demonstrating the model’s ability to adapt to in-context distributions.
Model  Shot Visual Question Answering Captioning
VQAv2 TextVQA OKVQA COCO NoCaps TextCaps
< 5B Model Comparisons
Flamingo-3B [22]  0 49.2 30.1 41.2 73.0
4 53.2 32.7 43.3 85.0
8 55.4 32.4 44.6 90.6
MM1-3B [9] 0 46.2 29.4 26.1 73.5 55.6 63.3
4 57.9 45.3 44.6 112.3 99.7 84.1
8 63.6 44.6 48.4 114.6 104.7 88.8
xGen-MM-base (4B) 0 43.1 34.0 28.0 67.2 82.6 69.5
4 66.3 54.2 48.9 107.6 100.8 89.9
8 66.9 55.3 50.1 109.8 104.6 94.0
Larger Models for Reference
Flamingo-9B [22] 8 58.0 33.6 50.0 99.0 - -
Idefics-9B [13] 8 56.4 27.5 47.7 97.0 86.8 63.2
MM1-7B [9] 8 63.6 46.3 51.4 116.3 106.6 88.2
Idefics2-8B [11] 8 70.3 57.9 54.6 116.0 - -
Table 1: Few-shot Pretraining Evaluation. Following [9], we randomly sample demonstrations from the training set as few-shot examples. We report CIDEr score for captioning and accuracy for VQA.
6.2 Supervised Fine-tuning
We evaluate our model on a comprehensive suite of multi-modal (image-text) benchmarks, assessing the model’s ability from multiple perspectives. Our evaluation covers general VQA benchmarks [74, 75, 76, 77, 78], visual perception [49], domain knowledge [79, 80], OCR ability [57, 81], and hallucination [82, 83]. For models fine-tuned on interleaved multi-image datasets, we also evaluate their performance on common multi-image benchmarks [69, 84, 85, 86].
Single-Image Evaluation.
In Table 2, we compare with models in comparable sizes ($<5$B), including both closed-source [9] and SoTA open-source models [10, 37]. We report individual benchmark scores along with two average scores: “Avg.(all)" is simply the average over all benchmarks, and “Avg.(perc.)" is the average score over benchmarks that focus on general VQA and visual perceptions. xGen-MM-instruct outperforms previous baselines on both general VQA and visual perception benchmarks. In addition, we find that xGen-MM-instruct-interleave, although further fine-tuned on multi-image data, maintains good performance on single-image benchmarks and has the highest overall scores.
Model (Size)
SEED
-IMG
SEED
v2
MMB
MM
Star
MME
CVB
-2D
CVB
-3D
RealW
QA
MMMU
Math
Vista
Sci
QA
POPE
Text
VQA
Avg.
Avg.
Closed-source models
GPT-4V 72.0 - 80.8 49.7 63.3 64.3 73.8 56.5 53.8 48.2 82.1 75.4 - - -
[…]
Open-source models
HPT-1.5-edge (4B) 72.3 - 74.6 45.8 - - - - 42.6 45.1 85.4 91.0 - - -
VILA-1.5-3B (3B) 67.9 - 63.4 - - - - - 33.3 - 69.0 85.9 - - -
VILA-1.5-3B (3B) 67.9 51.9 62.4 40.3 58.5 50.1 60.3 53.3 34.1 30.6 68.9 86.9 58.1 55.6 59.1
phi-3-vision (4B) - - 80.5 - - - - - - 44.5 90.8 85.8 70.9 - -
[…]
xGen-MM-inst.
-interleave (4B)
[…]
Table 2: Evaluation on single-image benchmarks. phi-3-vision and VILA-1.5-3B are tested with our evaluation code [^3] https://github.com/open-compass/VLMEvalKit for a fair comparison. We also include the GPT-4V (gpt-4-1106-preview) performance (provided by the evaluation codebase) as a reference in the first row.
Multi-Image Evaluation.
In Table 3, we compare xGen-MM-instruct with xGen-MM-instruct-interleave on multi-image benchmarks. Although the former is fine-tuned from xGen-MM-base which can comprehend interleaved image-text data, it performs poorly on multi-image benchmarks. We suspect it is because solely fine-tuning on single-image data hurts such ability. With multi-image SFT, we see significantly improved scores. In addition, we also evaluate Xgen-MM-interleave on single-image benchmarks (See Table 2) and find that it maintains good performance on all benchmarks with the highest overall scores.
Model BLINK QBench-2 Mantis-eval
GPT-4V 51.1 73.4 62.7
VILA-1.5-3B (3B) 39.8 51.7 41.9
xGen-MM-inst. (4B) 46.6 52.4 42.4
xGen-MM-inst.-interleave (4B) 49.7 75.1 56.7
Table 3: Evaluation on multi-image benchmarks. VILA-1.5-3B results are obtained using the same evaluation code as our models. We include the GPT-4V performance as a reference in the first row.
Figure 6: Example model outputs of xGen-MM-instruct-interleave. The model is capable of understanding interleaved image-text input and user queries about multiple images while maintaining the performance on single-image QAs.
6.3 Post-training
Table 4 summarizes the results of two post-training strategies for xGen-MM-instruct. We measure safety performance by ASR(attack success rate) on the VLGuard test split and hallucination performance using HallusionBench [82] (accuracy on image-context reasoning) and POPE [83] (average F1 score on binary entity presence questions). To ensure post-training doesn’t compromise helpfulness, we report performance on a few comprehension benchmarks as a control.
DPO enhances truthfulness by improving hallucination benchmarks (Row 2), while safety finetuning significantly reduces ASR (Row 3). Helpfulness is also improved slightly, as shown by control benchmarks. The final model, xGen-MM-dpo, includes both improvements.
Method Safety Hallucination Helpfulness (Control)
VLGuard ($\downarrow$) HalBench ($\uparrow$) POPE ($\uparrow$) SEED-IMG ($\uparrow$) MMB-dev ($\uparrow$) MME ($\uparrow$) MMStar ($\uparrow$)
xGen-MM-inst. (4B) 56.6 56.3 87.0 71.8 76.0 63.8 46.7
+ DPO 54.9 57.1 87.0 71.9 76.4 63.0 47.1
      + Safety FT 5.2 56.6 86.8 72.1 76.4 64.4 47.1
Table 4: Post-training results. We report results on safety and hallucination benchmarks after post-training, as well as on four helpfulness benchmarks as a control. Post-training improves harmlessness without compromising helpfulness.
7 Ablation Studies
7.1 Pre-training Ablation
Scaling Pre-training Data.
We perform an ablation study to explore the relation between the amount of pre-training data and the pre-train evaluation metrics, by varying the data scale from 2B multimodal tokens to 100B multimodal tokens. The data recipe we used here is a mixture of image caption datasets and multimodal interleaved data. As shown in Figure 7, we find that scaling up the number of multimodal tokens from 2B to 60B leads to substantial gain for image-text (COCO-Caps) and OCR (Text-Caps, TextVQA) tasks, and further increasing the data size to 100B has moderate additional benefit in terms of few-shot evaluation metrics.
(a) COCO-Caps
(b) Text-Caps
(c) OK-VQA
Figure 7: Few-shot performance given different sizes of pretraining data.
Pre-training Data Recipe.
We discuss the impact of different data recipes for pre-training. Specifically, we perform ablation studies on top of a base data recipe: use Obelics [13] as the multimodal interleaved data source while keeping the caption datasets mixture the same. We also consider two other recipes (1) using MINT-1T [12] as interleaved data replacement, and (2) mixing additional pure text-only instruction-tuning data as a pre-train dataset. As shown in Table 5, we see a performance improvement using MINT-1T for image-text alignment (COCO-Caps) and OCR (Text-Caps, TextVQA), with a slight performance drop on OK-VQA, which is a knowledge-intensive task. We also find that adding text data can help attain the performance on OK-VQA that relies more on LLM capacity.
Data Text-VQA OK-VQA COCO-Caps Text-Caps
Obelics 41.1 / 41.9 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
MINT-1T 41.0 / 42.0 46.5 / 48.2 109.3 / 111.4 80.3 / 82.0
MINT-1T + text data 42.1 / 42.6 48.3 / 49.7 108.0 / 110.2 77.0 / 79.9
Table 5: Few-shot (4-shot / 8-shot) performance given different data recipes.
Visual Backbones.
We also explore if different visual backbones have an impact on the performance of vision-language tasks. We compare two types of visual encoders, DFN and SigLIP. Empirically, we find SigLIP provides better visual representations that boost performance on OCR tasks.
Visual Backbone Text-VQA OK-VQA COCO-Caps Text-Caps
DFN 41.1 / 41.8 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
SigLIP 49.1 / 50.5 48.4 / 48.9 108.7 / 110.2 84.7 / 88.6
Table 6: Few-shot (4-shot / 8-shot) performance given different visual backbones.
Number of Visual Tokens.
Another ablation is to study the impact of different numbers of visual tokens, i.e., input image tokens fed into the language model. We find that reducing the number of visual tokens from 128 to 64 can still attain similar performance, as shown in Table 7. This makes it possible for models to take in more visual images given a fixed context window.
Visual Token Text-VQA OK-VQA COCO-Caps Text-Caps
128 41.1 / 41.8 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
64 41.2 / 42.6 47.6 / 48.3 108.0 / 109.3 79.5 / 81.6
Table 7: Few-shot (4-shot / 8-shot) performance given the different number of visual tokens.
7.2 SFT Ablation
We conduct ablation studies at the instruction fine-tuning stage, focusing on several model design choices and data recipes. The SFT ablation studies are conducted on a simplified SFT data mixture, so the results in this section are not directly comparable to the main results in section  6.2.
Any-Resolution Vision Token Sampling.
Our any-resolution strategy differs from previous work [46] in that every group of image embeddings (of the same image patch) is downsampled with a perceiver resampler, which ensures that the number of vision tokens input to the LLM remains relatively small. In this section, we ablate the effectiveness of our any-resolution strategy by comparing it with a “fixed-resolution" baseline and other downsampling designs.
The “fixed-resolution" baseline resizes all images to the default input size of the vision encoder while keeping the original aspect ratios. We also tried another downsampling strategy with the perceiver resampler: Instead of doing downsampling for each patch independently, we consider a "fixed sampling" (denoted as anyres-fixed-sampling in Figure 8(a)). In the fixed sampling, we concatenate the image embeddings from all image patches and then input them as a single sequence to the perceiver resampler to obtain the fixed number of vision tokens for the whole image.
Figure 8: SFT ablation studies. (a). Comparison of different vision token sampling strategies on OCR benchmarks. (b). Comparison between our model and its “instruction-aware" alternative. For each evaluation domain in Figure (b), we report the average score on multiple relevant benchmarks.
Our evaluation of this design focuses on text-rich tasks (e.g., document understanding) that would benefit from high-resolution encoding with visual details. From Figure 8(a), we can see significant improvements with our resolution image encoding strategy even with downsampled vision tokens. The fixed sampling strategy, although it shows improvements over the base resolution baseline, is not as good as the patch-wise sampling. We suspect that this may be due to two reasons: (a) With fixed sampling, a vision token sequence that can have as long as over 3,000 embedding tokens will be compressed to 128 tokens, which may be too few to retain the information. (b) The perceiver resampler may not work well with a concatenation of different image embeddings.
Instruction-Aware Vision Token Sampling.
InstructBLIP [7] proposes an instruction-aware Q-Former [1] for vision token sampling and shows that it can improve the model performance on some benchmarks. With the perceiver resampler as the VL connector, we can adopt a similar modification to make this process instruction-aware. To make our perceiver resampler “instruction-aware", we append the text instructions tokens to the query tokens of the perceiver resampler. Unlike Q-Former, there are only cross-attention layers inside the perceiver resampler, so the instruction (text tokens) would interact with both query tokens and image embeddings via cross-attention.
From the comparison in Figure 8(b), we do not observe a significant difference between our model and its instruction-aware version on various benchmarks. It could be that our modification to the perceiver resampler can not be identical to the instruction-aware modification made to the Q-Former in  Dai et al. [7], and thus the effectiveness differs. Because of the little difference we observe in this ablation study, we keep the original perceiver resampler architecture in our model for simplicity. We leave the further exploration of the “instruction-aware" VL connector to future works.
Text-only SFT data
MMMU
MathVista
Science
QA
MME
MMStar
Conversation 39.1 37.1 84.8 64.9 46.1
Conversation + Math + Coding 40.9 38.9 81.4 64.8 45.3
Table 8: The impact of text-only SFT data. We compare two choices of text-only SFT data used for the image-text SFT data mixture.
Quality of the Text-only Instruction Data.
It is a common strategy to train or fine-tune a multi-modal LLM on both multi-modal and pure text data [8, 11]. It is mainly for maintaining the language ability of the fine-tuned model. In this experiment, we study how this pure text subset would affect the performance on multi-modal benchmarks. For the instruction tuning stage, we compare whether the diversity of the pure text data would affect the multi-modal performance. For example, how pure-text math data affects a model’s performance on multimodal math benchmarks. In our main experiments, the default collection of pure text instruction data covers diverse domains including conversation [67], math [68, 87], and code [88]. For this ablation study, we substitute these datasets with the same amount of samples that only cover general conversation.
In Table 8, we observe that adding math and coding data, although in pure-text format, can help improve a model on relevant benchmarks like MathVista [80], while has less effects on general VQA benchmarks.
8 Conclusion
We introduce xGen-MM (BLIP-3), a comprehensive framework for training a series of open-source large multimodal models on a curated mixture of large-scale datasets. xGen-MM (BLIP-3) demonstrates emergent abilities such as multimodal in-context learning and achieves impressive results on multimodal benchmarks. By open-sourcing xGen-MM (BLIP-3), our curated datasets, and our SFT fine-tuning codebase, we hope to empower the research community with accessible multimodal foundation models and datasets, allowing practitioners to explore further and advance the potential and emergent abilities of LMMs.
9 Broader Impact
The xGen-MM (BLIP-3) framework and its suite of Large Multimodal Models (LMMs) have the potential to significantly advance multimodal AI research by providing accessible, open-source resources for the broader community. By facilitating the development and fine-tuning of state-of-the-art LMMs, xGen-MM (BLIP-3) empowers researchers and practitioners across various domains to innovate and apply these models to diverse real-world challenges. Moreover, the integration of safety-tuning protocols within the xGen-MM (BLIP-3) framework helps mitigate ethical risks such as bias and misinformation, promoting the responsible deployment of AI technologies.
10 Acknowledgement
We would like to thank Srinath Meadusani, Lavanya Karanam, Dhaval Dilip Metrani, and Eric Hu for their work on the scientific computation infrastructure, as well as Jason Lee and John Emmons for their efforts in collecting the large-scale text-only SFT datasets used in one of our pre-training ablation studies.

### v3

@@
-xGen-MM (BLIP-3): A Family of Open Large Multimodal Models
+BLIP-3: A Family of Open Large Multimodal Models
 Abstract
-This report introduces xGen-MM (also known as BLIP-3), a framework for developing Large Multimodal Models (LMMs).
+This paper introduces BLIP-3, an open framework for developing Large Multimodal Models (LMMs).
 The framework comprises meticulously curated datasets, a training recipe, model architectures, and a resulting suite of LMMs.
-xGen-MM, short for xGen-MultiModal, expands the Salesforce xGen initiative on foundation AI models.
+We release 4B and 14B models, including both the pre-trained base model and the instruction fine-tuned ones.
 Our models undergo rigorous evaluation across a range of tasks, including both single and multi-image benchmarks.
-Our pre-trained base model exhibits strong in-context learning capabilities and the instruction-tuned model demonstrates competitive performance among open-source LMMs with similar model sizes.
-In addition, we introduce a safety-tuned model with DPO, aiming to mitigate harmful behaviors such as hallucinations and improve safety.
-We open-source our models, curated large-scale datasets, and our fine-tuning codebase to facilitate further advancements in LMM research.
-Associated resources will be available on our project page above.
-Refer to caption Refer to caption (a) BLIP-2 framework (b) xGen-MM (BLIP-3) framework Figure 1: We introduce xGen-MM (BLIP-3), a framework (b) for developing Large Multimodal Models (LMMs).
+Our models demonstrate competitive performance among open-source LMMs with similar model sizes.
+Our resulting LMMs demonstrate competitive performance among open-source LMMs with similar model sizes, with the ability to comprehend interleaved image-text inputs.
+Our training code, models, and all datasets used in this work, including the three large-scale datasets we create and the preprocessed ones, will be open-sourced to better support the research community.
+Refer to caption Figure 1: We introduce BLIP-3, a framework (b) for developing Large Multimodal Models (LMMs).
 Our framework improves upon BLIP-2 (a) [42] by (1) increasing the richness, scale, and diversity of training data, (2) replacing the Q-Former layers with a more scalable vision token sampler, and (3) simplifying the training process via the unification of the training objectives to a single loss at every training stage.
-The resulting suite of LMMs can perform various visual language tasks and achieve competitive performance across benchmarks.
+(c) illustrates the stage-by-stage training process, from initializing with a pre-trained LLM to the final LMMs.
 1 Introduction
 Large Multimodal Models (LMMs) have attracted significant attention with their potential applications and emergent capabilities.
-Recent advancements in both proprietary models [2, 3, 4, 5] and open-source LMMs [6, 1, 7, 8, 9, 10, 11] highlight the rapid progress and growing interest in this field.
-However, despite these advancements, there is still a gap between open-source models and proprietary ones in terms of access to open weights, training recipes, and curated datasets.
-Such limitations hinder the open-source communities from replicating, understanding, and improving LMMs.
-Recent works have demonstrated that large-scale and high-quality data are essential for training robust LMMs [8, 9, 10, 11, 12].
+Recent advancements in both proprietary models [63, 25, 64, 82] and open-source LMMs [41, 42, 15, 49, 60, 47, 38, 29, 46, 2, 26, 55, 71, 66] highlight the rapid progress and growing interest in this field.
+However, despite these advancements, a significant gap remains between open-source and proprietary models, particularly in terms of accessibility to good base models, training recipes, and curated datasets.
+While there are many strong open-weight models [1, 2, 26], not many of them disclose their training details or release their full training data recipe (including training code).
+This lack of transparency limits the broader research community’s ability to replicate, analyze, and improve LMMs.
+Additionally, many of these models do not contribute new datasets, further restricting progress.
+These constraints hinder transparency and innovation, preventing open-source communities from fully leveraging and advancing LMM research.
+Recent works have demonstrated that large-scale and high-quality data are essential for training robust LMMs [49, 60, 47, 38, 6, 7].
 BLIP-2 [42] was one of the pioneering efforts in exploring LMMs, which leveraged synthetic data to achieve impressive results at the time (Figure 1 (a)).
@@
 Moreover, BLIP-2 supports only single-image input, whereas interleaved multimodal data formats are the most natural form of multimodal data [37].
-In response to these challenges, we introduce xGen-MM (BLIP-3) (Figure 1 (b)), a new framework designed to scale up LMM training by utilizing an ensemble of multimodal interleaved datasets, curated caption datasets, and other publicly available datasets [14, 15, 16, 17].
-xGen-MM, short for xGen-MultiModal, further expands our previous generative AI initiatives and corresponding foundation models for text xGen [18], code generation codeGen [19, 20], function calling APIGen [21], among others.
-In xGen-MM (BLIP-3), as illustrated in Figure 2, we streamline the model architecture by replacing the Q-Former [1] with a more scalable vision token sampler (specifically, a perceiver resampler [22]) and simplifying the training objectives to focus solely on the auto-regressive loss of text tokens in a multimodal context.
+In response to these challenges, we introduce BLIP-3 (Figure 1 (b)), a new framework designed to scale up LMM training by utilizing an ensemble of multimodal interleaved datasets, curated caption datasets, and other publicly available datasets [11, 36, 65, 22, 9].
+In BLIP-3, as illustrated in Figure 2, we streamline the model architecture by replacing the Q-Former [42] with a more scalable vision token sampler [4] and simplifying the training objectives to focus solely on the auto-regressive loss of text tokens in a multimodal context.
 Our primary focus is on dataset curation and scaling up the training data.
-Recently, our BLIP-3 team introduced two large-scale, high-quality datasets: MINT-1T [12], a trillion-token scale interleaved dataset; and BLIP3-KALE, a knowledge-augmented high-quality dense captions dataset.
-In this technical report, we introduce two additional specialized datasets: BLIP3-OCR-200M, a large-scale dataset with dense OCR annotations; and BLIP3-GROUNDING-50M, a large-scale visual grounding dataset.
-In addition to these datasets, we are committed to open-sourcing the series of models developed in this work, including the base, instruction-tuned, and DPO models.
-Along with the model release, we also provide code for easy fine-tuning of our base model on custom datasets.
-By making these resources publicly available, we aim to make LMM research and development more accessible to the community, and we encourage researchers and practitioners to use our models and datasets to understand and further explore the potential and emergent abilities of LMMs.
-Figure 2: Overview of the xGen-MM (BLIP-3) framework.
+Recently, several high-quality, large-scale datasets have been introduced, such as MINT-1T[6], a trillion-token-scale interleaved dataset, and BLIP3-KALE[7], a knowledge-augmented dataset with high-quality dense captions.
+In this paper, we introduce three additional specialized datasets: BLIP3-OCR-200M, a large-scale dataset with dense OCR annotations for general base-resolution OCR understanding; BLIP3-GROUNDING-50M, a large-scale visual grounding dataset; and BLIP3-OCR-HD-30M, a high-resolution dataset with high-quality OCR annotations, designed for high-resolution pre-training.
+In addition to these datasets, we are committed to open-sourcing the series of models developed in this work, including both the pre-trained base models and instruction-tuned models.
+Along with the model release, we also provide our training code.
+By making these resources publicly available, we aim to make LMM research and development more accessible to the community, and we encourage researchers and practitioners to use our models and datasets to understand and further explore the potential and emergent capabilities of LMMs.
+Figure 2: Overview of the BLIP-3 architecture.
 Free-form interleaved images and texts from the ensembled interleaved and caption datasets are input into the framework, with each modality undergoing a separate tokenization process to be fed into the pre-trained LLM in natural order.
 A standard auto-regressive loss is then applied to the text tokens.
-The Vision Transformer is kept frozen during training, while all other parameters, including the token sampler and the pre-trained LLM, are trained.
+The Vision Transformer is kept frozen during training, while all other parameters, including the vision token sampler and the pre-trained LLM, are trained.
 2 Related Work
-Recent advancements in Large Multimodal Models (LMMs) have explored two main architectural approaches: the cross-attention style [22, 23] and the decoder-only style.
-The cross-attention approach, exemplified by models like Flamingo [22, 23] and Llama 3.1 [5], integrates vision and language modalities through a complex attention mechanism to enable deep multimodal understanding.
-Another approach is the decoder-only architecture [1, 7, 8, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], which we adopt in xGen-MM (BLIP-3), offers a more streamlined solution.
-This approach connects pre-trained language models to visual inputs using lightweight connectors, simplifying the integration process while maintaining robust multimodal capabilities.
-The effectiveness of this architecture is evident in models such as MM1 [9], VILA [10], LLaVA [8], phi3-vision [37], and Otter [38].
+Recent advancements in Large Multimodal Models (LMMs) have explored different architectural designs.
+We roughly divide them into two main categories: the cross-attention style [4, 5] and the self-attention style.
+The cross-attention approach, exemplified by models like Flamingo [4, 5] and Llama 3.1 [55], integrates vision and language modalities by injecting visual information into layers inside LLMs using a cross-modal cross-attention module.
+The self-attention approach [42, 15, 49, 8, 14, 17, 30, 67, 48, 77, 78, 24, 10, 23, 70], on the other hand, offers a more streamlined solution.
+This approach connects pre-trained language models to visual inputs using lightweight connectors.
+Both visual and textual inputs are tokenized and concatenated in the same input sequence to LLM, and the LLM learns to align both modalities through self-attention.
+The self-attention design has become prevalent recently, which has been shown to be effective on a wide range of LMMs [50, 1, 2, 76, 8] .
+For BLIP-3, we also adopt the self-attention design due to its simplicity.
 Training methodologies for LMMs typically follow one of the two strategies.
@@
 The second strategy involves extensive pre-training on large-scale, diverse datasets, followed by visual instruction fine-tuning.
-This approach, employed by models like MM1 and Idefics2 [11], infuses broad knowledge into the model, which is then fine-tuned to align with human-like interactions and safety standards.
-While MM1 [9] provides extensive ablations and studies on the recipes aimed at improving LMMs, it releases limited resources for practitioners to adopt the model (e.g., MM1 models and datasets are close-sourced).
-Idefics2 [11] is a more recent open-source work that open-sources a suite of models along with detailed training strategies and data recipes, but Idefics2 mostly resues existing datasets in their experiments without contributing new ones.
-In this work, we present xGen-MM (BLIP-3).
-Unlike previous works, xGen-MM (BLIP-3) is an open-source family of a series of models, data recipes, fine-tuning code, and two large-scale foundational multimodal datasets, which we hope will enable and advance future research in this area.
+This approach, used by models such as MM1 series [60, 82] and Idefics2 [38], infuses a broad knowledge into the model during the pre-training stage, and then fine-tune it for better instruction following ability and better alignment.
+While MM1 series [60, 82] provides extensive ablations and studies on the recipes aimed at improving LMMs, it releases limited resources for practitioners to reproduce the model (their models and datasets are not open-sourced).
+In this work, we introduce BLIP-3, an open LMM family encompassing a series of models, data recipes, training code, and three new large-scale foundational multimodal datasets.
+Unlike previous works, BLIP-3 is designed to enable and advance future research in this area by providing comprehensive resources for the community.
 3 Model Architecture
 Architecture Overview.
-As illustrated in Figure 2, the xGen-MM (BLIP-3) framework adopts an architecture consisting of a ViT [44, 45], a vision token sampler (perceiver resampler [22]) to downsample the image embeddings, and a pre-trained Large Language Model (phi3-mini [37]).
+As illustrated in Figure 2, the BLIP-3 framework adopts an architecture consisting of a ViT [19, 81], a vision token sampler (perceiver resampler [4]) to downsample the image embeddings, and a pre-trained LLM (Phi-3 series [1]).
 The input to the model can be free-form multimodal interleaved texts and vision tokens from the diverse multimodal data sources we ensemble.
 Any-Resolution Vision Token Sampling.
-As proved effective in recent LMMs [46, 47, 48], we adopt a dynamic high-resolution (i.e., "any-resolution") image encoding strategy at the fine-tuning and post-training stages.
+As proved effective in recent LMMs [50, 83, 16], we adopt a dynamic high-resolution (i.e., “any-resolution”) image encoding strategy at both the Stage-2 pre-training and the fine-tuning stages.
 We enable higher-resolution image understanding with image patch-wise encoding.
@@
 4 Training
-Pre-training.
-The pre-training objective is to predict the next text token across the dataset mixture we pre-train on.
-Overall, the resulting base model xGen-MM-Phi3-mini-base-r is pre-trained for about 100 billion multimodal tokens from the ensembled dataset, and our pre-training resolution is 384x384 pixels, which aligns with SigLIP [45].
-Supervised Fine-tuning (SFT).
-We further fine-tune our pre-trained models on instruction-following examples to make them better understand and follow user queries.
-At the fine-tuning stage, we use a collection of publically available instruction-following datasets [11, 29, 49].
-We adopt the any-resolution vision token sampling strategy to allow a better understanding of images of higher resolutions such as text-rich document data.
-We introduce the technical details for the fine-tuning stage in the following sections.
+Stage-1 Base Resolution Pre-training.
+The first stage of pre-training focuses on base resolution image-text alignment using a diverse dataset mixture comprising open-source datasets and two newly created datasets from BLIP-3 – BLIP3-OCR-200M and BLIP3-GROUNDING-50M, all of which will be made publicly available.
+In this stage, the base model is trained on approximately 100 billion multimodal tokens from the ensembled dataset.
+To ensure computational efficiency, the pre-training resolution is set to 384×384 pixels, aligning with the default resolution of SigLIP [81].
+Stage-2 High Resolution Pre-training.
+Building upon the pre-trained checkpoint from Stage 1, the second stage incorporates additional high-quality, high-resolution datasets, including one newly created dataset from BLIP-3 – BLIP3-OCR-HD-30M, to further enhance the model’s capability.
+To maintain consistency with the supervised fine-tuning (SFT) stage, we adopt the same any-resolution vision token sampling strategy introduced in section 3.
+We set the any-resolution grid to support up to 12 image patches with a maximum side length of 1536 pixels on the longer side.
+Further details and ablation studies are provided in section 7.2.
+Single-Image Supervised Fine-tuning.
+Then, we fine-tune the pre-trained model on a collection of publicly available instruction-following datasets [38, 48, 71].
+During fine-tuning, we employ the same any-resolution vision token sampling strategy as in Stage 2, enabling the model to process high-resolution images effectively, including text-rich document-style data.
+The following sections provide additional technical details on the fine-tuning procedure.
 Interleaved Multi-Image Supervised Fine-tuning.
-We conduct a second-stage fine-tuning on the instruction fine-tuned model on a mixture of multi-image and single-image instructions-following samples.
-The goal for this second-stage fine-tuning is to enhance the model’s ability to comprehend interleaved image-text input, which is helpful for multimodal in-context learning, multi-image question answering, and many more practical use cases.
-For the multi-image fine-tuning, we also adopt the any-resolution vision token sampling strategy same as in the previous SFT stage.
-Post-training.
-Finally, we perform two stages of post-training to improve the model’s helpfulness while mitigating harmful qualities such as hallucination and toxicity.
-We first perform direct preference optimization (DPO [50]) to improve the model’s helpfulness and visual faithfulness.
-We then perform safety fine-tuning to improve the model’s harmlessness.
-We quantitatively demonstrate Pareto gains in model harmlessness and helpfulness after our post-training.
+We conduct a second-stage fine-tuning on the instruction fine-tuned model on a mixture of multi-image and single-image instruction-following samples.
+The goal of this second-stage fine-tuning is to enhance the model’s ability to comprehend interleaved image-text input, which is helpful for multimodal in-context learning, multi-image question answering, and many more practical use cases.
+For the multi-image fine-tuning, we also adopt the any-resolution vision token sampling strategy, the same as in the previous SFT stage.
 5 Data
 5.1 Pre-training Data Recipe
-As indicated in Figure 3, in xGen-MM (BLIP-3), we pre-train on an ensemble of diverse multimodal datasets with the indicated sampling ratios.
-Figure 3: Overview of xGen-MM (BLIP-3) Pre-training Datasets.
+Our Stage-1 pre-training recipe is detailed in Figure 3. In BLIP-3, we pre-train on a diverse set of multimodal datasets using the specified sampling ratios.
+For Stage-2 pre-training, we use the Stage-1 data recipe as a base while incorporating specialized high-resolution, text-rich datasets.
+Specifically, we replace BLIP3-OCR-200M with BLIP3-OCR-HD-30M and introduce the IDL [9] dataset, each making up 10% of the total training data.
+To preserve the overall structure of the dataset mixture, we adjust the sampling ratios of the remaining caption datasets by reducing BLIP3-KALE from 25% to 10%, Datacomp-1B from 10% to 2.5%, and BLIP3-GROUNDING-50M from 5% to 2.5%, while keeping the rest unchanged.
+Figure 3: Overview of BLIP-3 Stage-1 pre-training datasets base recipe.
 Interleaved Dataset Mixture.
 We combine MINT-1T (including its HTML, PDF, and ArXiv subsets) with OBELICS (HTML only) to create a more diverse and comprehensive dataset mixture that covers a broader range of domains.
-  1. 1.
-    MINT-1T [12] is a trillion token scale multimodal interleaved dataset, containing data sources from HTML, PDF, and ArXiv.
+1. MINT-1T [6].
+A trillion-token-scale multimodal interleaved dataset, containing data sources from HTML, PDF, and ArXiv.
 As evidenced by MM1 [60] and Idefics2 [38], such multimodal interleaved datasets are essential for scaling up large multimodal model training and enabling fundamental capabilities like multimodal in-context learning.
-Notably, different from the OBELICS [11], MINT-1T has three subsets from different sources: the HTML subset, the PDF subset, and the ArXiv subset.
-In xGen-MM (BLIP-3), these three subsets are mixed in a 7:5:1 ratio.
-  2. 2.
-    OBELICS [11] is another large-scale multimodal interleaved dataset constructed from HTML documents solely.
-It differs slightly in domain coverage from MINT-1T due to the specific preprocessing steps adopted.
-Caption Dataset Mixture.
-We integrate a diverse range of caption datasets, with the specifics outlined in the following details.
-  1. 1.
-    BLIP3-KALE is a large-scale curated high-quality caption dataset. Details will be discussed in another paper, and the dataset will be made public very soon.
-  2. 2.
-    BLIP3-OCR-200M is a curated large-scale OCR dataset to address the limitations of current large multimodal models in handling text-rich images like documents and charts, as traditional image-text datasets often lack adequate OCR annotations.
-To enhance text comprehension abilities, we use the OCR engine PaddleOCR [51] to annotate images with OCR-specific annotations.
-Overall, we curate a dataset of 200 million high-resolution images from Datacomp-1B[17].
-For each image, we create captions with OCR data by identifying and extracting textual elements using the off-the-shelf OCR engine [51].
-Text segments in a caption like "… text …" are modified to include OCR information as "… text ( ocr_info ) …", where ocr_info contains bounding box coordinates for the extracted text, specifying its exact position within the image in the format "<bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>".
+In BLIP-3 we mix its HTML, PDF, and ArXiv subsets in a 7:5:1 ratio.
+2. OBELICS [38].
+A multimodal interleaved dataset constructed solely from HTML documents. It differs slightly in domain coverage from MINT-1T due to the specific preprocessing steps adopted.
+Caption Dataset Mixture. We integrate a diverse range of caption datasets, with the following details.
+Figure 4: Samples from BLIP3-OCR-200M.
+We extract six levels of OCR granularity, with and without bounding boxes.
+OCR-related captions in BLIP-3 are preprocessed to remove filler phrases like ’the text’ within the dataloader, which we find improving OCR benchmarks’ performance.
+Samples from BLIP3-OCR-HD-30M.
+Structured similarly to BLIP3-OCR-200M, this dataset contains high-resolution images annotated with high-resolution OCR tools for enhanced high resolution OCR understanding.
+Only Level 1 is shown for simplicity.
+Samples from BLIP3-GROUNDING-50M.
+A large-scale dataset of images and corresponding captions containing localization information about objects.
+Furthermore, we release the associated object bounding box data to facilitate the creation of captions with custom templates.
+1. BLIP3-KALE [7].
+An open-source large-scale curated high-quality caption dataset, which we use as a base for general detailed caption datasets.
+2. BLIP3-OCR-200M.
+A curated large-scale OCR dataset with 200 million images from Datacomp- 1B[22].
+For each image, we use the off-the-shelf OCR engine [28] for OCR annotation.
+Text segments in a caption like "… text …" are modified to include OCR information as "… text ( ocr_info ) …", where ocr_info contains normalized bounding box coordinates for the extracted text, specifying its exact position within the image in the format "<bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>".
 We have multiple granularities of OCR information, including with and without bounding box data.
-In our work, we only utilize textual information without bounding box data, which has proven to be effective.
-Note that, in xGen-MM (BLIP-3), we preprocess the captions to remove filler texts such as “the text”, which we find improves OCR-related benchmark performance.
-We hypothesize that this is because such filler text is relatively easy to predict, potentially diluting the loss associated with OCR-relevant tokens.
-Nonetheless, incorporating bounding box information could further enhance performance, and we encourage researchers in the community to explore this potential.
-  3. 3.
-    BLIP3-GROUNDING-50M is a curated large-scale grounding dataset to enhance the ability to ground semantic concepts in visual features, which is crucial for tasks like object detection, semantic segmentation, and understanding referring expressions [52] (e.g., "the object to the left of the dog").
+In BLIP-3 pre-training, we only utilize textual information without bounding box data.
+3. BLIP3-GROUNDING-50M.
+A curated large-scale grounding dataset to enhance the ability to ground semantic concepts in visual features, which is crucial for tasks like understanding referring expressions [79] (e.g., "the object to the left of the dog").
 We curate a dataset of 50 million images from Datacomp-1B [22].
@@
 Objects mentioned in a caption like "… object …" are modified to include grounding information as " … object ( grounding_info ) …", where grounding_info contains bounding box information in one of three formats, each capturing a different granularity of localization: (1) <bbox>$x_{1},y_{1},x_{2},y_{2}$</bbox>, (2) "starts at $(x_{1},y_{1})$ and extends up to $(x_{2},y_{2})$", or (3) "top-left corner of the image".
-  4. 4.
-    Other Public Datasets Mixture: We also include other publicly available datasets such as uncurated Datacomp-1B [17] image-text pairs, CC12M [14], CC3M [14], VG [15], and SBU [16].
-Figure 4: Samples from BLIP3-OCR-200M.
-Six levels of OCR information granularity are extracted, with and without bounding box data.
-Note that in xGen-MM (BLIP-3), OCR-related captions are preprocessed to remove filler phrases like ’the text,’ resulting in improved OCR benchmark performance.
-Figure 5: Samples from BLIP3-GROUNDING-50M.
-We introduce a large-scale dataset of images and corresponding captions containing localization information about objects.
-Furthermore, we release the associated object bounding box data to facilitate the creation of captions with custom templates.
+4. BLIP3-OCR-HD-30M.
+A high-resolution OCR dataset curated for Stage-2 pre-training.
+Different from BLIP3-OCR-200M, which focuses on more general OCR data with fewer resolution constraints, this dataset is filtered to only contain images of which both the height and width are larger than 512 pixels.
+On the selected high-resolution images, we run PaddleOCR with a high-resolution setup to obtain more accurate OCR annotations.
+Our empirical study shows that running the OCR annotation process with a high-res setting can help get a more complete and accurate set of OCR annotations that matches its resolution, which is more useful for the Stage-2 pre-training.
+This process yields a total of 30 million high-resolution samples with detailed OCR annotations that match the image quality.
+5.Other Datasets Mixture.
+We also include other publicly available datasets such as uncurated Datacomp-1B [22] image-text pairs, CC12M [11], CC3M [11], VG [36], SBU [65], and IDL [9].
 5.2 Supervised Fine-tuning Data Recipe
-The datasets used in the fine-tuning stage are from public datasets of different domains [29, 11, 49].
-We sample data with various domain focuses including multi-modal conversation [29], image captioning [55, 56], visual question answering [57, 58, 59, 60], chart/document understanding [61, 62, 63, 64], science and math [65, 66].
-In addition to the multi-modal image-text data, we also mix in pure text instruction following data  [67, 68] during visual instruction tuning.
-Ultimately, we collect a mixture of 1 million publically available instruction-tuning samples, on which we fine-tune our model for one epoch.
-The multi-image instruction tuning stage starts with a model fine-tuned on single-image samples.
+The datasets used in the fine-tuning stage are from public datasets of different domains [48, 38, 71, 75, 12, 69, 61, 68, 31, 59, 58, 34, 35, 33, 56, 84].
+In addition to the multi-modal image-text data, we also mix in pure text instruction-following data  [45, 62] during visual instruction tuning.
+Ultimately, we collect a mixture of 3 million publicly available instruction-tuning samples, on which we fine-tune our model for one epoch.
+The multi-image instruction tuning stage starts with a model fine-tuned on single-image data.
 We use a mixture of public multi-image / interleaved image-text instruction data [32, 54].
 To prevent the model from deteriorating on single-image capabilities, we reuse a subset of single-image datasets used in the previous fine-tuning stage and mix them into the multi-image training data.
-5.3 Post-training Data Recipe
-Improving Truthfulness by Direct Preference Optimization.
-We employ VLFeedback [71], a synthetically annotated multimodal preference dataset that uses off-the-shelf VLMs to generate responses to a diverse mix of multimodal instructions that are then scored by GPT4-V [2] along three axes – helpfulness, visual faithfulness, and ethics.
-The dataset contains 80k such instructions from which we construct preference data by marking as preferred (and dispreferred) the response with the highest (and lowest) average score across models and filtering out examples with low-scoring preferred responses.
-We thus generate 62.6k preference examples.
-We perform 1 epoch of DPO on the combined preference dataset while updating a subset (2.5%) of LLM backbone weights using low-rank adaptation (LoRA [72]).
-Also, following recent work [50], we generate an additional set of responses that capture the model’s intrinsic hallucinations, by performing a second step of DPO per-iteration against the models’ output to a noised version of the input image and original query, which we treat as an additional dispreferred response.
-Improving Harmlessness by Safety Fine-tuning.
-Next, we perform 3 epochs of safety fine-tuning on the train split of the VLGuard [73] dataset, which contains 2k examples of unsafe images and instructions.
-VLGuard comprises two types of unsafe examples: (1) objectionable images paired with safe instructions and a desirable abstention response, and (2) safe images paired with two types of instruction-response pairs, one safe and another unsafe.
-The dataset consists of unsafe examples belonging to various subcategories including privacy-violating, risky/sensitive topics (such as politics, sex, and violence), deception, and discrimination.
-Following the original work, we randomly sample 5k additional examples from the instruction fine-tuning dataset to retain the model’s helpfulness without exaggerating its safety behavior.
-As before, we update a subset (2.5%) of LLM backbone weights using low-rank adaptation.
 6 Experiments
-6.1 Pre-training
-Few-shot Evaluation.
-After the pre-training stage, we evaluate our pre-trained model on classic captioning and VQA tasks, in comparison with previous models that support few-shot learning multi-modal evaluation.
-We present zero-shot and few-shot (4- and 8-shots) results, as shown in Table 1.
-Overall, our model achieves competitive multimodal in-context learning performance with comparable-sized LMMs  [^1] For few-shot evaluation, the zero-shot results are used mainly as a reference for their corresponding few-shot numbers, and they can be sensitive to prompts.
-As also mentioned in [9], they are mostly indicative of how well the pre-training distribution matches the associated evaluation task format.
-In pre-training evaluation, we mainly care for few-shot performance, which is robust to prompt templates..
-For the OCR tasks (TextCaps and TextVQA) and VQA-v2, it significantly outperforms MM1-3B and even larger models such as Idefics-9B [13] and MM1-7B [9].
-On all benchmarks, increasing the number of shots can improve the performance, demonstrating the model’s ability to adapt to in-context distributions.
-Model  Shot Visual Question Answering Captioning
-VQAv2 TextVQA OKVQA COCO NoCaps TextCaps
-< 5B Model Comparisons
-Flamingo-3B [22]  0 49.2 30.1 41.2 73.0
-4 53.2 32.7 43.3 85.0
-8 55.4 32.4 44.6 90.6
-MM1-3B [9] 0 46.2 29.4 26.1 73.5 55.6 63.3
-4 57.9 45.3 44.6 112.3 99.7 84.1
-8 63.6 44.6 48.4 114.6 104.7 88.8
-xGen-MM-base (4B) 0 43.1 34.0 28.0 67.2 82.6 69.5
-4 66.3 54.2 48.9 107.6 100.8 89.9
-8 66.9 55.3 50.1 109.8 104.6 94.0
-Larger Models for Reference
-Flamingo-9B [22] 8 58.0 33.6 50.0 99.0 - -
-Idefics-9B [13] 8 56.4 27.5 47.7 97.0 86.8 63.2
-MM1-7B [9] 8 63.6 46.3 51.4 116.3 106.6 88.2
-Idefics2-8B [11] 8 70.3 57.9 54.6 116.0 - -
-Table 1: Few-shot Pretraining Evaluation. Following [9], we randomly sample demonstrations from the training set as few-shot examples. We report CIDEr score for captioning and accuracy for VQA.
-6.2 Supervised Fine-tuning
-We evaluate our model on a comprehensive suite of multi-modal (image-text) benchmarks, assessing the model’s ability from multiple perspectives.
-Our evaluation covers general VQA benchmarks [74, 75, 76, 77, 78], visual perception [49], domain knowledge [79, 80], OCR ability [57, 81], and hallucination [82, 83].
+Model Accessibility
+SEED-IMG
+MMB(dev)
+MMStar
+MME(norm)
+RWQA
+MMVet
+MMMU(val)
+MathVista
+TextVQA
+OCRBench
+POPE
+HalBench
+Average
+open-
+weight
+open-
+data
+open-
+code
+data
+contribution
+GPT-4V [63] $\times$ $\times$ $\times$ $\times$ 72.0 80.8 49.7 63.3 61.4 49.0 53.8 48.2 78.0 67.8 81.8 39.3 62.1
+3B-5B Models
+MM1.5-3B [82] $\times$ $\times$ $\times$ $\times$ 72.4 - - 64.2 56.9 41.0 37.1 44.4 76.5 65.7 88.1 - -
+Phi-3-vision [1] $\times$ $\times$ $\times$ 71.0 73.6 47.9 55.3 58.8 43.2 46.1 45.1 73.3 63.7 83.5 39.0 58.4
+Phi-3.5-vision [1] $\times$ $\times$ $\times$ 69.6 81.9 47.3 65.6 53.5 43.2 43.0 43.9 72.0 59.9 86.1 39.6 58.8
+MiniCPM-V-2.0 [76] $\times$ $\times$ $\times$ 67.1 69.1 39.1 64.6 55.8 41.0 38.2 39.8 73.2 59.6 86.3 36.1 55.8
+VILA-1.5-3B [47] $\times$ 67.9 62.4 40.3 58.5 53.3 38.5 34.1 30.6 58.1 43.7 86.9 31.2 50.5
+BLIP-3-4B (ours) 71.9 74.7 45.3 62.5 61.0 41.2 42.3 44.5 76.5 66.0 87.9 36.5 59.2
+10B-15B Models
+Pixtral-12B [2] $\times$ $\times$ $\times$ 71.5 77.9 54.5 68.6 65.4 58.5 51.1 56.3 75.7 72.2 84.2 47.0 64.9
+Llama-3.2-11B-V-inst. [55] $\times$ $\times$ $\times$ 72.7 68.0 49.8 65.0 63.3 57.6 48.0 47.7 54.1 75.3 88.1 40.3 60.8
+VILA-1.5-13B [47] $\times$ 68.0 74.4 44.2 61.4 53.3 45.0 41.1 42.3 61.1 45.7 85.0 39.3 55.1
+Cambrian-13B [71] 73.2 73.2 47.1 67.0 58.6 48.9 41.6 47.7 72.8 61.0 86.8 39.4 59.8
+BLIP-3-14B (ours) 73.4 77.5 48.7 68.1 64.4 44.5 47.0 45.4 79.0 71.1 86.0 38.7 62.0
+Table 1: Evaluation on single-image benchmarks.
+For benchmarks where a open-source model doesn’t report its official score, we use the score reported in a third-party leaderboard [^2] https://huggingface.co/spaces/opencompass/open_vlm_leaderboardor run it ourselves using the evaluation codebase [18] for a fair comparison.
+We also include the GPT-4V (gpt-4-1106-preview) performance (provided by the evaluation codebase) as a reference in the first row.
+We evaluate our models (4B and 14B) on a comprehensive suite of multimodal benchmarks, assessing the model’s ability from multiple perspectives.
+Our evaluation covers general VQA benchmarks [39, 52, 20, 13, 3], domain knowledge [80, 57], OCR ability [69, 53], and hallucination [27, 44].
 For models fine-tuned on interleaved multi-image datasets, we also evaluate their performance on common multi-image benchmarks [32, 74, 72, 21].
 Single-Image Evaluation.
-In Table 2, we compare with models in comparable sizes ($<5$B), including both closed-source [9] and SoTA open-source models [10, 37].
-We report individual benchmark scores along with two average scores: “Avg.(all)" is simply the average over all benchmarks, and “Avg.(perc.)" is the average score over benchmarks that focus on general VQA and visual perceptions.
-xGen-MM-instruct outperforms previous baselines on both general VQA and visual perception benchmarks.
-In addition, we find that xGen-MM-instruct-interleave, although further fine-tuned on multi-image data, maintains good performance on single-image benchmarks and has the highest overall scores.
-Model (Size)
-SEED
--IMG
-SEED
-v2
-MMB
-MM
-Star
-MME
-CVB
--2D
-CVB
--3D
-RealW
-QA
-MMMU
-Math
-Vista
-Sci
-QA
-POPE
-Text
-VQA
-Avg.
-Avg.
-Closed-source models
-GPT-4V 72.0 - 80.8 49.7 63.3 64.3 73.8 56.5 53.8 48.2 82.1 75.4 - - -
-[…]
-Open-source models
-HPT-1.5-edge (4B) 72.3 - 74.6 45.8 - - - - 42.6 45.1 85.4 91.0 - - -
-VILA-1.5-3B (3B) 67.9 - 63.4 - - - - - 33.3 - 69.0 85.9 - - -
-VILA-1.5-3B (3B) 67.9 51.9 62.4 40.3 58.5 50.1 60.3 53.3 34.1 30.6 68.9 86.9 58.1 55.6 59.1
-phi-3-vision (4B) - - 80.5 - - - - - - 44.5 90.8 85.8 70.9 - -
-[…]
-[…]
-xGen-MM-inst.
--interleave (4B)
-[…]
-Table 2: Evaluation on single-image benchmarks.
-phi-3-vision and VILA-1.5-3B are tested with our evaluation code [^3] https://github.com/open-compass/VLMEvalKit for a fair comparison.
-We also include the GPT-4V (gpt-4-1106-preview) performance (provided by the evaluation codebase) as a reference in the first row.
+In Table 1, we compare models of comparable sizes, including both closed-source [63, 82] and open-source models [47, 1, 76, 2, 55, 71].
+We report individual benchmark scores as well as the overall average score across all benchmarks, following the standard practice.
+Model BLINK QBench MuirBench Mantis-Eval
+GPT-4V [63] 51.1 73.4 - 62.7
+3B-5B Models
+VILA-1.5-3B [47] 39.8 51.7 29.2 41.9
+BLIP-3-4B-SI 42.9 51.9 35.0 49.3
+BLIP-3-4B-MI 47.1 69.6 38.0 56.2
+10B - 15B Models
+VILA-1.5-13B [47] 48.1 61.2 37.7 49.3
+BLIP-3-14B-SI 47.1 55.4 49.1 55.3
+BLIP-3-14B-MI 53.9 73.4 56.2 61.3
+Table 2: Evaluation on multi-image benchmarks. VILA-1.5 models are evaluated using the same evaluation code as our models.
 Multi-Image Evaluation.
-In Table 3, we compare xGen-MM-instruct with xGen-MM-instruct-interleave on multi-image benchmarks.
-Although the former is fine-tuned from xGen-MM-base which can comprehend interleaved image-text data, it performs poorly on multi-image benchmarks.
-We suspect it is because solely fine-tuning on single-image data hurts such ability.
-With multi-image SFT, we see significantly improved scores.
-In addition, we also evaluate Xgen-MM-interleave on single-image benchmarks (See Table 2) and find that it maintains good performance on all benchmarks with the highest overall scores.
-Model BLINK QBench-2 Mantis-eval
-GPT-4V 51.1 73.4 62.7
-VILA-1.5-3B (3B) 39.8 51.7 41.9
-xGen-MM-inst. (4B) 46.6 52.4 42.4
-xGen-MM-inst.-interleave (4B) 49.7 75.1 56.7
-Table 3: Evaluation on multi-image benchmarks. VILA-1.5-3B results are obtained using the same evaluation code as our models. We include the GPT-4V performance as a reference in the first row.
-Figure 6: Example model outputs of xGen-MM-instruct-interleave.
-The model is capable of understanding interleaved image-text input and user queries about multiple images while maintaining the performance on single-image QAs.
-6.3 Post-training
-Table 4 summarizes the results of two post-training strategies for xGen-MM-instruct.
-We measure safety performance by ASR(attack success rate) on the VLGuard test split and hallucination performance using HallusionBench [82] (accuracy on image-context reasoning) and POPE [83] (average F1 score on binary entity presence questions).
-To ensure post-training doesn’t compromise helpfulness, we report performance on a few comprehension benchmarks as a control.
-DPO enhances truthfulness by improving hallucination benchmarks (Row 2), while safety finetuning significantly reduces ASR (Row 3).
-Helpfulness is also improved slightly, as shown by control benchmarks.
-The final model, xGen-MM-dpo, includes both improvements.
-Method Safety Hallucination Helpfulness (Control)
-VLGuard ($\downarrow$) HalBench ($\uparrow$) POPE ($\uparrow$) SEED-IMG ($\uparrow$) MMB-dev ($\uparrow$) MME ($\uparrow$) MMStar ($\uparrow$)
-xGen-MM-inst. (4B) 56.6 56.3 87.0 71.8 76.0 63.8 46.7
-+ DPO 54.9 57.1 87.0 71.9 76.4 63.0 47.1
-      + Safety FT 5.2 56.6 86.8 72.1 76.4 64.4 47.1
-Table 4: Post-training results.
-We report results on safety and hallucination benchmarks after post-training, as well as on four helpfulness benchmarks as a control.
-Post-training improves harmlessness without compromising helpfulness.
+In Table 2, we compare BLIP-3 single-image (SI) model with BLIP-3 interleaved multi-image (MI) model on multi-image benchmarks.
+Although the former is fine-tuned from a pre-trained model that can comprehend interleaved image-text data, it performs poorly on multi-image benchmarks, possibly because the single-image SFT causes degradation in such ability.
+With interleaved multi-image SFT, we see significant improvements.
 7 Ablation Studies
-7.1 Pre-training Ablation
+7.1 Ablation on Stage-1 Pre-training
+Few-shot Pre-training Evaluation.
+To analyze the impact of pre-training configurations without the impact of fine-tuning data, we conduct ablation studies on pre-trained models with few-shot evaluation.
+Following [60], we randomly sample a few-shot subset from the training set and evaluate model performance.
+We report CIDEr scores for captioning tasks and accuracy for VQA to provide a clear comparison of few-shot performance across different pre-training configurations.
 Scaling Pre-training Data.
-We perform an ablation study to explore the relation between the amount of pre-training data and the pre-train evaluation metrics, by varying the data scale from 2B multimodal tokens to 100B multimodal tokens.
-The data recipe we used here is a mixture of image caption datasets and multimodal interleaved data.
-As shown in Figure 7, we find that scaling up the number of multimodal tokens from 2B to 60B leads to substantial gain for image-text (COCO-Caps) and OCR (Text-Caps, TextVQA) tasks, and further increasing the data size to 100B has moderate additional benefit in terms of few-shot evaluation metrics.
+For Stage-1 pre-training, we perform an ablation study to explore the relation between the amount of pre-training data and the pre-train few-shot evaluation metrics, by varying the data scale from 2B multimodal tokens to 100B multimodal tokens.
+The data recipe we used here is a mixture of image caption datasets and multimodal interleaved data as detailed in 3.
+As shown in Figure 5, we find that scaling up the number of multimodal tokens from 2B to 60B leads to substantial gain for image-text (COCO-Caps) and OCR (Text-Caps) tasks, and further increasing the data size to 100B has moderate additional benefit in terms of few-shot evaluation metrics.
 (a) COCO-Caps
 (b) Text-Caps
-(c) OK-VQA
-Figure 7: Few-shot performance given different sizes of pretraining data.
-Pre-training Data Recipe.
-We discuss the impact of different data recipes for pre-training.
-Specifically, we perform ablation studies on top of a base data recipe: use Obelics [13] as the multimodal interleaved data source while keeping the caption datasets mixture the same.
-We also consider two other recipes (1) using MINT-1T [12] as interleaved data replacement, and (2) mixing additional pure text-only instruction-tuning data as a pre-train dataset.
-As shown in Table 5, we see a performance improvement using MINT-1T for image-text alignment (COCO-Caps) and OCR (Text-Caps, TextVQA), with a slight performance drop on OK-VQA, which is a knowledge-intensive task.
-We also find that adding text data can help attain the performance on OK-VQA that relies more on LLM capacity.
-Data Text-VQA OK-VQA COCO-Caps Text-Caps
-Obelics 41.1 / 41.9 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
-MINT-1T 41.0 / 42.0 46.5 / 48.2 109.3 / 111.4 80.3 / 82.0
-MINT-1T + text data 42.1 / 42.6 48.3 / 49.7 108.0 / 110.2 77.0 / 79.9
-Table 5: Few-shot (4-shot / 8-shot) performance given different data recipes.
+Figure 5: Few-shot performance given different scales of pre-training data in Stage-1 pre-training.
 Visual Backbones.
 We also explore if different visual backbones have an impact on the performance of vision-language tasks.
-We compare two types of visual encoders, DFN and SigLIP.
-Empirically, we find SigLIP provides better visual representations that boost performance on OCR tasks.
+We compare two types of visual encoders, DFN [19] and SigLIP [81].
+As shown in 3, we find SigLIP provides better visual representations that boost performance on OCR tasks, and we adopt SigLIP in the final model architecture as the ViT backbone.
+To ensure computational efficiency and a fair comparison, all models in this ablation study are pre-trained with 10B tokens.
 Visual Backbone Text-VQA OK-VQA COCO-Caps Text-Caps
-DFN 41.1 / 41.8 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
+DFN 41.1 / 41.9 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
 SigLIP 49.1 / 50.5 48.4 / 48.9 108.7 / 110.2 84.7 / 88.6
-Table 6: Few-shot (4-shot / 8-shot) performance given different visual backbones.
-Number of Visual Tokens.
-Another ablation is to study the impact of different numbers of visual tokens, i.e., input image tokens fed into the language model.
-We find that reducing the number of visual tokens from 128 to 64 can still attain similar performance, as shown in Table 7.
-This makes it possible for models to take in more visual images given a fixed context window.
-Visual Token Text-VQA OK-VQA COCO-Caps Text-Caps
-128 41.1 / 41.8 48.4 / 49.5 107.2 / 109.4 78.2 / 79.9
-64 41.2 / 42.6 47.6 / 48.3 108.0 / 109.3 79.5 / 81.6
-Table 7: Few-shot (4-shot / 8-shot) performance given the different number of visual tokens.
-7.2 SFT Ablation
-We conduct ablation studies at the instruction fine-tuning stage, focusing on several model design choices and data recipes.
-The SFT ablation studies are conducted on a simplified SFT data mixture, so the results in this section are not directly comparable to the main results in section  6.2.
+Table 3: Few-shot (4-shot / 8-shot) performances for different visual backbones at Stage-1 pre-training.
+7.2 Ablation on Stage-2 Pre-training
+In this section, we investigate the impact of Stage-2 pre-training.
+Model performance is evaluated by fine-tuning on a smaller subset ($\sim$ 1M samples) of our instruction tuning dataset for better computational efficiency.
+For evaluation, we focus on two representative LMM capabilities: general VQA and OCR, each of which is an averaged score over multiple relevant benchmarks.
+Impact of Adding Stage-2 Pre-training and with Different Resolutions.
+As shown in Table 4, we use our 100B token Stage-1 pre-trained model as the baseline.
+We then add Stage-2 pre-training on top of it; we use half the data samples for a full Stage-2 pre-training run to maintain computational efficiency.
+Furthermore, we examine how varying Stage-2 pre-training resolutions (i.e., different numbers of image patches) affect performance.
+To ensure a fair comparison, all models in Table 4 are fine-tuned with a maximum of 13 image patches and follow the same fine-tuning protocol.
+From the results, we observe that incorporating Stage-2 pre-training leads to a significant improvement in OCR, while general VQA performance remains stable.
+Stage-2 Stage-2 Maximum Resolutions OCR GeneralVQA
+$\times$ N/A 64.6 63.5
+(768x768), (384x1152), (1152x384) 65.3 63.2
+(1152x1536), (1536x1152) 67.2 63.3
+Table 4: Stage-2 pre-training and its resolution impact.
+We ablate two resolution grid settings for Stage-2 pre-training: one with up to 4 image patches, at most arranged as (2×2, 1×3, or 3×1), and another with up to 12 image patches, at most arranged as (3×4 or 4×3), with each patch sized at 384×384 pixels.
+Figure 6: Example model outputs of BLIP-3-4B-MI (interleaved multi-image model).
+The model is capable of understanding interleaved image-text input and user queries about multiple images while maintaining the performance on single-image QAs.
+7.3 Ablation on Instruction Tuning
+The ablation study in this subsection focuses on several model design choices.
+Evaluations are conducted on single-image fine-tuned checkpoints.
+Similar to section 7.2, models in this ablation are fine-tuned on a smaller SFT subset for better computational efficiency, so the results are not directly comparable to our main results.
+Perceiver Resampler vs.
+MLP.
+We experiment with two designs of vision token sampler with similar “vision compression ratios”.
+Both models are trained from Stage-1 pre-training.
+For perceiver resample, we use 128 query tokens per image patch to achieve a compression ratio of $r_{c}=729/128=5.7$ With MLP projector, to get a similar compression ratio, we append a 2x2 max pooling layer to it to have $r_{c}=4$.
+In Table 5, we compare the two models on benchmarks focusing on various domains and report the average score on each domain.
+Perceiver resampler, with a slightly higher compression ratio, has a lower score on OCR, but performs better in other domains, and has a better overall score.
+We therefore choose to use perceiver resampler for BLIP-3, because the higher compression ratio is crucial for efficiency when handling interleaved multi-image input.
+Vision Token Sampler GeneralVQA OCR Sci.& Math Hallucination
+Perceiver resampler 63.1 70.4 43.4 62.2
+MLP + 2x2 pooling 61.7 72.2 41.8 62.7
+Table 5: Ablation study on vision token sampler. We experiment with two sampling designs with similar vision compression ratio.
 Any-Resolution Vision Token Sampling.
-Our any-resolution strategy differs from previous work [46] in that every group of image embeddings (of the same image patch) is downsampled with a perceiver resampler, which ensures that the number of vision tokens input to the LLM remains relatively small.
-In this section, we ablate the effectiveness of our any-resolution strategy by comparing it with a “fixed-resolution" baseline and other downsampling designs.
-The “fixed-resolution" baseline resizes all images to the default input size of the vision encoder while keeping the original aspect ratios.
-We also tried another downsampling strategy with the perceiver resampler: Instead of doing downsampling for each patch independently, we consider a "fixed sampling" (denoted as anyres-fixed-sampling in Figure 8(a)).
-In the fixed sampling, we concatenate the image embeddings from all image patches and then input them as a single sequence to the perceiver resampler to obtain the fixed number of vision tokens for the whole image.
+Our any-resolution strategy differs from previous work [50] in that every group of image embeddings (of the same image patch) is downsampled with a perceiver resampler to ensure efficiency.
+In this section, we ablate the effectiveness of our any-resolution strategy by comparing it with other designs.
+Models in this study are fine-tuned from the Stage-1 pre-trained base resolution models.
 Figure 7: SFT ablation studies.
@@
 (b).
-Comparison between our model and its “instruction-aware" alternative.
+Comparison between our model and its “instruction-aware” alternative.
 For each evaluation domain in Figure (b), we report the average score on multiple relevant benchmarks.
-Our evaluation of this design focuses on text-rich tasks (e.g., document understanding) that would benefit from high-resolution encoding with visual details.
-From Figure 8(a), we can see significant improvements with our resolution image encoding strategy even with downsampled vision tokens.
+The “fixed-resolution” baseline resizes all images to the default input size of the vision encoder while keeping the original aspect ratios.
+Another downsampling strategy with the perceiver resampler is “fixed sampling” ( anyres-fixed-sampling), for which we concatenate the image embeddings from all image patches and then input them as a single sequence to the perceiver resampler to obtain the fixed number of vision tokens for the whole image.
+Our evaluation of this design focuses on text-rich OCR benchmarks.
+From Figure 7(a), we see significant improvements with the any-res image encoding strategy even with downsampled vision tokens.
 The fixed sampling strategy, although it shows improvements over the base resolution baseline, is not as good as the patch-wise sampling.
-We suspect that this may be due to two reasons: (a) With fixed sampling, a vision token sequence that can have as long as over 3,000 embedding tokens will be compressed to 128 tokens, which may be too few to retain the information.
+We suspect two reasons for this: (a) With fixed sampling, the vision token compression ratio can be too high to retain text-rich visual information.
 (b) The perceiver resampler may not work well with a concatenation of different image embeddings.
 Instruction-Aware Vision Token Sampling.
-InstructBLIP [7] proposes an instruction-aware Q-Former [1] for vision token sampling and shows that it can improve the model performance on some benchmarks.
-With the perceiver resampler as the VL connector, we can adopt a similar modification to make this process instruction-aware.
-To make our perceiver resampler “instruction-aware", we append the text instructions tokens to the query tokens of the perceiver resampler.
-Unlike Q-Former, there are only cross-attention layers inside the perceiver resampler, so the instruction (text tokens) would interact with both query tokens and image embeddings via cross-attention.
-From the comparison in Figure 8(b), we do not observe a significant difference between our model and its instruction-aware version on various benchmarks.
-It could be that our modification to the perceiver resampler can not be identical to the instruction-aware modification made to the Q-Former in  Dai et al.
-[7], and thus the effectiveness differs.
+InstructBLIP [15] proposes an instruction-aware Q-Former [42] for vision token sampling and proves its effectiveness on certain benchmarks.
+With the perceiver resampler, we experiment with a similar strategy by appending text instruction tokens to the query tokens of the perceiver resampler.
+This enables the instruction (text tokens) to interact with both query tokens and image embeddings via cross-attention.
+From the comparison in Figure 7(b), we do not observe a significant difference between our model and its instruction-aware version on common task domains.
 Because of the little difference we observe in this ablation study, we keep the original perceiver resampler architecture in our model for simplicity.
-We leave the further exploration of the “instruction-aware" VL connector to future works.
-Text-only SFT data
-MMMU
-MathVista
-Science
-QA
-MME
-MMStar
-Conversation 39.1 37.1 84.8 64.9 46.1
-Conversation + Math + Coding 40.9 38.9 81.4 64.8 45.3
-Table 8: The impact of text-only SFT data. We compare two choices of text-only SFT data used for the image-text SFT data mixture.
-Quality of the Text-only Instruction Data.
-It is a common strategy to train or fine-tune a multi-modal LLM on both multi-modal and pure text data [8, 11].
-It is mainly for maintaining the language ability of the fine-tuned model.
-In this experiment, we study how this pure text subset would affect the performance on multi-modal benchmarks.
-For the instruction tuning stage, we compare whether the diversity of the pure text data would affect the multi-modal performance.
-For example, how pure-text math data affects a model’s performance on multimodal math benchmarks.
-In our main experiments, the default collection of pure text instruction data covers diverse domains including conversation [67], math [68, 87], and code [88].
-For this ablation study, we substitute these datasets with the same amount of samples that only cover general conversation.
-In Table 8, we observe that adding math and coding data, although in pure-text format, can help improve a model on relevant benchmarks like MathVista [80], while has less effects on general VQA benchmarks.
+We leave the further exploration of the “instruction-aware” VL connector to future works.
 8 Conclusion
-We introduce xGen-MM (BLIP-3), a comprehensive framework for training a series of open-source large multimodal models on a curated mixture of large-scale datasets.
-xGen-MM (BLIP-3) demonstrates emergent abilities such as multimodal in-context learning and achieves impressive results on multimodal benchmarks.
-By open-sourcing xGen-MM (BLIP-3), our curated datasets, and our SFT fine-tuning codebase, we hope to empower the research community with accessible multimodal foundation models and datasets, allowing practitioners to explore further and advance the potential and emergent abilities of LMMs.
-9 Broader Impact
-The xGen-MM (BLIP-3) framework and its suite of Large Multimodal Models (LMMs) have the potential to significantly advance multimodal AI research by providing accessible, open-source resources for the broader community.
-By facilitating the development and fine-tuning of state-of-the-art LMMs, xGen-MM (BLIP-3) empowers researchers and practitioners across various domains to innovate and apply these models to diverse real-world challenges.
-Moreover, the integration of safety-tuning protocols within the xGen-MM (BLIP-3) framework helps mitigate ethical risks such as bias and misinformation, promoting the responsible deployment of AI technologies.
-10 Acknowledgement
-We would like to thank Srinath Meadusani, Lavanya Karanam, Dhaval Dilip Metrani, and Eric Hu for their work on the scientific computation infrastructure, as well as Jason Lee and John Emmons for their efforts in collecting the large-scale text-only SFT datasets used in one of our pre-training ablation studies.
+We introduce BLIP-3, a comprehensive framework for training a series of open-source large multimodal models on a curated mixture of large-scale datasets.
+BLIP-3 LMMs achieve competitive results on a range of multimodal benchmarks.
+By open-sourcing BLIP-3 (4B and 14B), our curated datasets, and our training code, we hope to empower the research community with better accessible multimodal foundation models and datasets, allowing practitioners to explore further and advance the potential and emergent abilities of LMMs.